OpenStudy (anonymous):

In a soccer tournament, teams receive 6 points for a winning game, 3 points for a trying game, and 1 point for each goal they score. what algebraic expression models the total number of points that a soccer team receives in a tournament?????? suppose one team wins two games and ties 1 game, scoring a total of 5 goals. how many points does the team receive????

OpenStudy (anonymous):

let: x = games won y = games played z = goal p = points p = 6x + 3y + z ----- team won 2 games, ties 1 and scores total of 5 goals in all 3 games played. total point? p = 6(2) + 3(1) +(5) = 20 points

OpenStudy (anonymous):

wouldn't it be =5 instead of +5?? just wondering

OpenStudy (anonymous):

for my answer I got 12w+3T+5g=(20points) is that correct????

OpenStudy (anonymous):

what would be 5 or +5 ? do you mean the term for goals ? the total points are sum of each individual points for; games won; tried and goals i am confused by the term "tried" in the question, does a tied game and lost game both count as "tried" game ?

OpenStudy (anonymous):

it doesn't say tried in the question maybe I had a typo... 3 points for a tying game

OpenStudy (anonymous):

your answer is correct, that is total is 20 points, but the expression is not. see the coefficients in your equation: 12w+3T+5g=20 points when you substitute values in variables, variables have to be omitted. the equation has to be: total points --> p = 6W + 3T + G substituting the wins, ties and goals in the above expression. p = 6(2) + 3(1) + 5 = 20 <-- see variables, W, T and G were substituted with values given in the question